### v3

# Traversable wormholes inside anisotropic magnetized neutron stars: physical properties and potential observational imprints

Abstract
In this paper, we formulate wormhole-plus-neutron-star (WH+NS) systems supported by two scalar fields, allowing for both pressure anisotropy of the neutron fluid and magnetic field. In general, such WH+NS systems contain ghosts; however, these ghosts can be eliminated. We find that the wormhole remains traversable regardless of whether anisotropy of the neutron fluid and/or magnetic fields are included. In particular, the null energy condition (NEC) remains violated in the vicinity of the wormhole throat, ensuring the traversable nature of the geometry. For magnetized configurations, the resulting WH+NS systems can become extremely massive, with ADM masses exceeding $8\,M_{\odot}$, and can exhibit large surface redshifts exceeding $z\simeq 1.5$. Furthermore, we analyze the gravitational-wave echo time of the systems, which serves as a potential observational imprint. Our results indicate that the echo time can vary depending on the fluid anisotropy and the magnetic field configuration, suggesting that WH+NS systems may provide distinctive signals of gravitational echo.
journal:
1 Introduction
Wormholes are intriguing constructs that serve as tunnels connecting distant spacetime points Dai2020 ; Turimov2025 ; Bronnikov2023 ; Cataldo2017 ; Yousaf2025 ; Prat2015 ; Dai2019 ; Godani2019 . Both distinct spacetime domains are linked through a throat-like geometry. Originally, wormholes emerged as exact solutions of the Einstein field equation Zangeneh2025 . Due to their geometric nature, these exotic objects may potentially exist in curved spacetimes within a broad class of gravity theories Bambi2021 . Interestingly, wormholes have been associated with rapid interstellar travel, warp drives Shinkai2015 , and candidates for active galactic nuclei Piotrovich .
Historically, the first insight into what would later be interpreted as a wormhole can be traced back to Flamm’s 1916 work Flamm1916 (see Ref. Flamm2015 for the English translation), where he analyzed the Schwarzschild spacetime. In 1935, Einstein & Rosen Einstein1935 found a solution bearing a resemblance to a wormhole, which was identified as a “bridge” that connects two identical physical spaces. This work gave rise to the term “Einstein–Rosen bridge”. The term “wormhole” was later introduced in 1957 by Misner & Wheeler Misner1957 . It is worth noting that the putative wormhole in the Einstein–Rosen bridge is not traversable, since it is colocated with the blackhole’s singularity. The surface of the blackhole is the event horizon, which cannot be a wormhole Radhakrishnan2024 .
The formalism of traversable wormholes was introduced by Morris & Thorne Morris1988 in their seminal 1988 paper. In constructing traversable wormholes, the null energy condition (NEC) must be violated by the energy-momentum tensor (EMT) Lobo2016 . It should be noted that there are three other energy conditions: the weak energy condition (WEC), the strong energy condition (SEC), and the dominant energy condition (DEC) Radhakrishnan2024 . The violation of NEC implies that all the energy conditions are not validated Moraes2019 . In general relativity (GR), the violation of NEC typically requires the presence of exotic matter Kokubu2020 . Many authors Chianese2017 ; Mustafa2020 ; AzregAinou2015 ; Kuhfittig2006 ; Bhattacharya2019 ; Sreekumar2025 have included the exotic matter in their studies on wormholes. Furthermore, a comprehensive review of wormhole physics can be found in Ref. Lobo2016 . Refs. Falco1 ; Falco2 ; Falco3 ; Falco4 ; Falco5 ; Falco6 work on the development of possible astrophysical techniques to detect the wormholes.
Some authors have highlighted the potential existence of wormholes under extreme astrophysical conditions. Bhar et al. Bhar2016 investigated the possibility of sustaining static, spherically symmetric traversable wormhole geometries in GR, supported by dark energy and admitting conformal motion. Biswas et al. Biswas2024 proposed that wormholes could mimic supermassive black holes at galactic centers. Refs. Hao1 ; Hao2 work on wormhole-plus-black-hole systems. Other extreme astrophysical environments where wormholes are hypothesized to exist include compact stars Dzhunushaliev2011 ; Dzhunushaliev2012 ; Dzhunushaliev2013 ; Dzhunushaliev2014 ; Aringazin2015 ; Dzhunushaliev2015 ; Dzhunushaliev2016 ; Dzhunushaliev2023 ; Nojiri2024 . However, in Refs. Dzhunushaliev2011 ; Dzhunushaliev2012 ; Dzhunushaliev2013 ; Dzhunushaliev2014 ; Aringazin2015 ; Dzhunushaliev2015 ; Dzhunushaliev2016 ; Dzhunushaliev2023 , the existence of hybrid systems composed of wormholes and compact stars is sustained by the presence of ghosts, which leads to undesirable consequences: both mixed star-plus-wormhole systems and pure wormholes are unstable Nojiri2024 . Nojiri et al. Nojiri2024 eliminated the ghosts by imposing the constraints given by the additional Lagrange multipliers.
It has to be noted that although the model proposed in Ref. Nojiri2024 has successfully constructed a formalism of ghost-free systems, it did not take into account the anisotropy that may arise. It is widely known that the matter inside compact stars like neutron stars (NSs) may be anisotropic, meaning that the radial and tangential pressures differ. Various factors can lead to such pressure anisotropy; including boson condensation, the presence of a solid core, different types of phase transitions, strong magnetic or electric fields, and the influence of modified spacetime geometry Pattersons2021 . In line with these physical considerations, several studies on compact stars Pattersons2021 ; Sulaksono2014 ; Sulaksono2015 ; Silva2015 ; Setiawan2017 ; Lopes2019 ; Rizaldy2024 ; Posada2024 ; Becerra2024 ; Pattersons2025 ; Yusmantoro2025 ; Sakti2016 ; Becerra2024b ; Becerra2025 ; Torres2019 ; Rahmansyah2020 have incorporated anisotropic pressure. Anisotropy plays an important role in the mass calculation of compact stars. For instance, Ref. Pattersons2025 shows that anisotropy can enhance the mass of NSs, allowing them to reach the mass range of the secondary compact object in GW190814, which has been reported in Ref. Abbott2019 and hypothesized as the heaviest NS ever observed Fattoyev2020 . The review of anisotropic compact stars can be referred to Ref. Kumar2022 .
Another aspect that is not taken into account in the model presented in Ref. Nojiri2024 is the presence of magnetic fields. The magnetic fields on NSs are typically in the range of $10^{12}$$10^{15}$Lander2009 . There are classes of pulsars, such as anomalous X-ray pulsars (AXPs) and soft gamma repeater pulsars (SGRs), that have been identified as producing strong magnetic fields. SGRs are associated with supernova remnants, which correspond to young NSs Rizaldy2018a . Furthermore, observations of some AXPs also indicate that their surface magnetic fields are around $10^{14}-10^{15}$Rizaldy2018a . For further details on the observational diversity of magnetized NSs, see also Ref. Enoto2019 , while a general review of NSs is provided in Ref. Vidaña2018 .
One framework for describing magnetic fields in NSs is through chaotic magnetic fields, which has been investigated by Lopes & Menezes Lopes2015 . In their work, they also introduced an ansatz in which the magnetic field is coupled to the energy density of the star. An important advantage of adopting a chaotic magnetic field is the elimination of another anisotropy induced by the magnetic field, thereby simplifying the mathematical formulation. In addition to this simplification, the use of chaotic magnetic fields allows one to obtain a magnetic pressure that is independent of both direction and the choice of coordinate system, thus restoring the proper thermodynamic concept of pressure and ensuring consistency with field theory Lopes2015 ; Pattersons2024a . Several studies have subsequently incorporated chaotic magnetic fields in their analyses Pattersons2024a ; Wu2017 ; Lopes2020 ; Pattersons2024b .
Our aim in this work is to extend the wormhole plus neutron star (WH+NS) model constructed in Ref. Nojiri2024 by incorporating the anisotropy of the neutron fluid and the effects of chaotic magnetic fields that may arise in such a system. Other distinctions of our work from previously studied WH+NS configurations constructed in Refs. Dzhunushaliev2011 ; Dzhunushaliev2012 ; Dzhunushaliev2013 ; Dzhunushaliev2014 ; Aringazin2015 ; Dzhunushaliev2015 ; Dzhunushaliev2016 ; Dzhunushaliev2023 ; Nojiri2024 is the inclusion of the calculations of surface redshift and the echo time. Motivated by the possibility that WH+NS systems with an anisotropic neutron fluid and a chaotic magnetic field may contribute to gravitational-wave echoes characterized by the echo time, our results provide potential observational imprint for future LIGO/Virgo searches.
We emphasize that the presence of a magnetic field should not be regarded as a merely trivial extension of the system studied in Ref. Nojiri2024 . This is clearly reflected in the fact that the magnetized WH+NS configuration admits a different analytical solution for the throat radius. In addition, we obtain potential observational imprints of magnetized WH+NS systems that have not been explored previously, most notably their gravitational mass, which can exceed $8\,M_{\odot}$, being noticeably higher than that of the ordinary WH+NS systems investigated in Ref. Nojiri2024 , and their surface redshift, which can be larger than 1.5—both significantly higher than those of ordinary NSs. Moreover, we find that the combined effects of fluid anisotropy and the magnetic field lead to distinctive echo times, providing further observationally relevant signatures.
The paper is organized as follows. In Sect. 2, we introduce the formalisms of the anisotropic fluid and the chaotic magnetic field. Sec. 3 presents the formulation of GR coupled with two scalar fields, and demonstrates the elimination of ghosts. In this section, we also analyze the energy conditions and the travesability of the wormholes. In Sec. 4, we construct a model of non-magnetized WH+NS systems that consider anisotropy. Sec. 5 shows the model of anisotropic magnetized WH+NS systems. Note that, in both Secs. 4 and 5, we compare the obtained WH+NS systems with ordinary NSs; and also present the calculation results of the surface redshift and the echo time that may arise in such systems. Finally, Sec. 6 is devoted to the conclusions of our study.
2 Anisotropic fluid and chaotic magnetic field
To make this section self-contained, in Subsect. 2.1 we introduce the anisotropic fluid model employed in this study, while the formulation and ansatz of the chaotic magnetic field are given in Subsect. 2.2.
2.1 Model of anisotropy
It is important to consider the EMT of matter with anisotropic pressure, given by Pretel2023
$T_{\text{matter}\,\mu\nu}=(\rho+p_{t})u_{\mu}u_{\nu}+p_{t}\,g_{\mu\nu}+\sigma\,\xi_{\mu}\xi_{\nu},$ (1)
where $\rho$ is the energy density, $p_{t}$ is the tangential pressure, and $\sigma=p_{r}-p_{t}$ denotes the anisotropy term, where $p_{r}$ is the radial pressure. At this stage, we have to emphasize that $\rho$, $p_{r}$, and $p_{t}$ belong to matter. In Eq. (1), $u^{\mu}$ is the four-velocity of the fluid, satisfying the normalization condition $u^{\mu}u_{\mu}=-1$, and $\xi^{\mu}$ is a unit radial four-vector, satisfying $\xi^{\mu}\xi_{\mu}=1$. The vector $\xi_{\mu}$ is physically an auxiliary quantity that parametrizes the anisotropic stresses in the neutron fluid. Its components are fully determined by the geometry and matter profiles, in such a way so that $p_{r}=\xi^{\mu}\xi^{\nu}T_{\text{matter}\>\mu\nu}$, and $p_{t}=\varsigma^{\mu}\varsigma^{\nu}T_{\text{matter}\>\mu\nu}$, where $\varsigma^{\mu}$ is a unit spacelike vector (orthogonal to $u^{\mu}$ and $\xi^{\mu}$anisoadd .
By choosing $\xi_{\mu}=(0,\sqrt{g_{rr}},\>0,\>0)$, for static spherically symmetric configuration, we can easily find
$T_{\text{matter}\>\mu\nu}=\begin{pmatrix}-g_{tt}\>\rho&0&0&0\\ 0&g_{rr}\>p_{r}&0&0\\ 0&0&g_{\theta\theta}\>p_{t}&0\\ 0&0&0&g_{\varphi\varphi}\>p_{t}\end{pmatrix}$. (2)
In this work, we consider Cosenza-Herrera-Esculpi-Witten (CHEW) model Cosenza1981 in the realization of the WH+NS with anisotropic neutron fluid.
For the static spherically symmetric spacetime, the metric describing the interior of ordinary NSs is given by
${\color[rgb]{0,0,0}ds^{2}=-e^{2\nu(r)}dt^{2}+e^{2\lambda(r)}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),}$ (3)
where $\lambda(r)=\left(1-\frac{2m(r)}{r}\right)^{-1}$, and $m(r)$ denotes mass of the NSs as a function of $r$. Using the EMT shown in Eq. (1), Tolman-Oppenheimer-Volkoff (TOV) equation for anisotropic NSs reads
${\color[rgb]{0,0,0}\frac{dp_{r}}{dr}=-\frac{d\nu}{dr}(\rho+p_{r})-\frac{2\sigma}{r}.}$ (4)
Authors in Ref. Cosenza1981 considered that $\sigma/r$ can be written as
${\color[rgb]{0,0,0}\frac{\sigma}{r}=f(p_{r},r)(\rho+p_{r}),}$ (5)
where $f(p_{r},r)$ is an arbitrary function. Motivated to obtain a minimal modification of the TOV equation, authors in Ref. Cosenza1981 assumed $f(p_{r},r)=\frac{1-h}{2}\frac{d\nu}{dr}$, where $h$ is a constant. On the other words, it is not assumed for any specific physical reason but only because it transforms TOV equation into the simple form, i.e.
${\color[rgb]{0,0,0}\frac{dp_{r}}{dr}=-h\frac{d\nu}{dr}(\rho+p_{r})}.$ (6)
Finally, the expression of the anisotropy term can also be written as Rahmansyah2020
$\sigma=-\frac{r}{2}\frac{1-h}{h}\frac{dp_{r}}{dr},$ (7)
where $h$ denotes the free parameter of anisotropy. The system becomes isotropic when we set $h=1$.
2.2 Chaotic magnetic fields
[…]
$p_{\mathcal{B}}=\frac{1}{3}<T^{j}_{j}>\>=\frac{1}{3}\left(\frac{\mathcal{B}^{2}}{8\pi}+\frac{\mathcal{B}^{2}}{8\pi}-\frac{\mathcal{B}^{2}}{8\pi}\right).$ (8)
Here, $T^{j}_{j}$ is the spatial component of the EMT.
The magnetic energy density $\rho_{\mathcal{B}}$ and magnetic pressure $p_{\mathcal{B}}$ can now conveniently be written as
$\rho_{\mathcal{B}}=\frac{\mathcal{B}^{2}}{8\pi},\>\>\>\>\>p_{\mathcal{B}}=\frac{\mathcal{B}^{2}}{24\pi}.$ (9)
It should be noted that the standard treatment of magnetic fields in NSs, which leads to the appearance of anisotropy even though the original matter of the NSs is isotropic Bordbar2022 , complicates the mathematical formulation and prevents the stellar configuration from remaining spherically symmetric Mallick2014 .
According to Ref. Lopes2015 , the chaotic magnetic field approximation adopted here is physically motivated by the expectation that magnetic fields generated during the turbulent core-collapse supernova phase are highly disoriented. In this framework, the magnetic pressure is assumed to be statistically isotropic, so that no global preferred direction exists and anisotropic magnetic stresses associated with ordered field configurations are effectively averaged out. This allows the magnetic contribution to be incorporated consistently within the spherically symmetric formalism through an effective isotropic pressure. A limitation of the model is that it is expected that the magnetic field evolves and becomes at least partially oriented. The oriented magnetic field is a necessity to the theory of pulsars as celestial lighthouse. In this work, the the chaotic magnetic field is employed as an approximation to calculate the influence of the magnetic field in NS properties within the spherically symmetric TOV formalism.
Another important point to note is that the exact mathematical profile of the magnetic field inside NSs remains unknown. Lopes & Menezes Lopes2015 proposed an ansatz in which the magnetic field is coupled to the matter energy density $\rho$, i.e.
$\mathcal{B}=B_{s}+B_{0}\left(\frac{\rho}{\rho_{0}}\right)^{\eta},$ (10)
where $B_{s}$ denotes the magnetic field at the surface of the NS, $B_{0}$ represents the magnetic field expected at the core. For the WH+NS configuration, $B_{0}$ can be interpreted as the magnetic field expected at the wormhole’s throat. Moreover, $\rho_{0}$ denotes the central energy density of the non-magnetized NS configuration (without the wormhole) corresponding to the maximum mass, and $\eta$ is an arbitrary positive constant. It is worth noting that choosing $\eta$ to be very close to zero would make the magnetic field profile resemble a constant magnetic field. To simplify the calculation of the radius of the wormhole throat in Sec. 5, we set $\eta=1$ throughout this work. Moreover, the magnetic field parameters that are varied in our analysis are both $B_{0}$ and $B_{s}$.
It is important to note that the internal magnetic field profile of NSs is still subject to significant uncertainties. For this reason, it is customary to introduce phenomenological but physically motivated ansatz for the magnetic field profile. Different prescriptions have been proposed in the literature, including magnetic field profiles coupled to the baryon number density, as in Ref. Mallick2014 , as well as profiles coupled to the energy density.
In the latter approach, i.e. the ansatz proposed in Ref. Lopes2015 , the energy density can enter directly into the TOV equations governing the macroscopic structure of ordinary NSs. The parameter $\rho_{0}$ serves as a reference energy density scale, while the parameter $\eta$ controls the strength of the magnetic field. This choice provides a convenient and physically motivated parametrization for exploring the impact of strong magnetic fields on NS properties.
3 GR coupled with two scalar fields and a magnetic field
For completeness, Subsect. 3.1 contains the derivation of the field equations from the action, Subsect. 3.2 addresses the procedure for eliminating the ghosts, and Subsect. 3.3 analyzes the energy conditions and traversability of the wormholes.
3.1 Einstein field equations
The action of GR coupled with 2 scalar fields and magnetic field is given by
Here, $A_{\phi}$ is the derivative of $A(\phi,\chi)$ with respect to $\phi$, etc.
3.2 Elimination of the ghosts
First, it is important to note that the elimination of ghosts presented in this subsection follows the procedure outlined in Ref. Nojiri2024 . Now consider a metric of a general spherically symmetric and time-dependent spacetime, given by
$ds^{2}=$ $-e^{2\nu(t,r)}dt^{2}+e^{2\lambda(t,r)}dr^{2}$
$+r^{2}(d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}).$ (15)
Let us now assume the ansatz
$\phi=t,\quad\chi=r.$ (16)
A detailed justification that this ansatz does not lead to any loss of generality can be found in Refs. Nojiri2021 ; NojiriNashed2024 ; Nojiri2024NPB .
According to Ref. Nojiri2024 ; Nojiri2021 ; NojiriNashed2024 ; Nojiri2024NPB ; Nojiri2023 , the functions $A(\phi,\chi)$ and/or $C(\phi,\chi)$ are often negative, which implies that $\phi$ and/or $\chi$ behave as ghosts. To eliminate the ghosts, appropriate constraints must be introduced by adding Lagrange multipliers $\Lambda_{\phi}$ and $\Lambda_{\chi}$. This results in the addition of an additional action $S_{\Lambda}$, which reads Nojiri2024 ; NojiriNashed2024
${\color[rgb]{0,0,0}S_{\Lambda}}=\int d^{4}x\sqrt{-g}\big[$ ${\color[rgb]{0,0,0}\Lambda_{\phi}}\big(e^{-2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1\big)$
$+{\color[rgb]{0,0,0}\Lambda_{\chi}}\big(e^{-2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1\big)\big].$ (17)
Variations of $S_{\Lambda}$ with respect to $\Lambda_{\phi}$ and $\Lambda_{x}$ give the following constraints:
$0$ $=e^{-2\nu(t=\phi,r=\chi)}\partial_{\mu}\phi\partial^{\mu}\phi+1,$ (18)
$0$ $=e^{-2\lambda(t=\phi,r=\chi)}\partial_{\mu}\chi\partial^{\mu}\chi-1.$ (19)
The fluctuations of $\phi$ and $\xi$ can be written as Nojiri2024 ; NojiriNashed2024 ; Nojiri2024NPB
$\phi=t+\delta\phi,\>\>\>\ \chi=r+\delta\chi.$ (20)
By using Eqs. (18) and (19), one can obtain
$\partial_{t}(\delta\phi)=\partial_{r}(\delta r)=0,$ (21)
which means that both $\delta\phi$ and $\delta\chi$ vanish in the whole spacetime if we set the initial condition $\delta\phi=0$ and the boundary condition $\delta\chi\rightarrow 0$ when $r\rightarrow\infty$, and we are left with $\delta\phi=\delta\chi=0$ Nojiri2024 ; NojiriNashed2024 ; Nojiri2024NPB .
It is worth noting that, as pointed out in Ref. Nojiri2024 , even within the model defined by the modified action $S_{GR\phi\chi\mathcal{B}}+{\color[rgb]{0,0,0}S_{\Lambda}}$, the choice $\color[rgb]{0,0,0}\Lambda_{\phi}=\Lambda_{\chi}=0$} consistently satisfies the field equations. Consequently, any solution of Eq. (LABEL:field) derived from the original action (LABEL:action) also remains a solution in the modified model with action $S_{GR\phi\chi\mathcal{B}}+{\color[rgb]{0,0,0}S_{\Lambda}}$.
It is also worth noting that the constraints given by Eqs. (18) and (19) generalize the mimetic constraint introduced in Ref. Chamseddine2013 , where nondynamical dark matter effectively emerges. The review of mimetic gravity can be referred to Ref. Vagnozzi . In the present model—consistent with the discussion in Ref. Nojiri2024 —what effectively arises is not nondynamical dark matter, but rather nondynamical exotic matter like a phantom.
It is important to clarify the role of the auxiliary scalar fields $\chi$ and $\phi$ introduced in this construction. These fields are not intended to represent independent propagating degrees of freedom, but rather serve as auxiliary variables that parametrize the geometry and generate an effective stress–energy tensor through the structure of the action. In particular, once the ansatz $\chi=r$ and $\phi=t$ is imposed, the functions $A(\phi,\chi)$, $C(\phi,\chi)$, and the potential $V(\phi,\chi)$ become fixed and give rise to an effective exotic matter sector capable of supporting the wormhole geometry.
From a physical perspective, the exotic matter emerges as an effective source induced by the constrained auxiliary fields. As the construction remains free of ghost instabilities, the scalar fields should therefore be understood as non-propagating auxiliary fields with a physical sourcing role, rather than as conventional matter fields or pure coordinate labels.
3.3 Energy conditions and traversability
First, we have to highlight that the total EMT in our model generally consist of contributions from ordinary matter, the magnetic field, and the scalar fields. In GR, the NEC, WEC, SEC, and DEC impose the following sequence of inequalities on the total EMT $T_{\text{tot}\,\mu\nu}$ Nojiri2024 ; Carroll2004 :
To obtain the traversable wormhole, we need the violation of the NEC in the vicinity of a throat Zaslavskii2007 . It implies the violation of all energy conditions Nojiri2024 . Before analyzing the traversability of the wormhole, we must first determine the total energy density $\rho_{\text{tot}}$, the total radial pressure $p_{r,\text{tot}}$, and the total tangential pressure $p_{t,\text{tot}}$. Once again we have to highlight that these quantities consist of contributions from ordinary matter, the magnetic field, and the scalar fields. To obtain them, one must first derive the explicit expressions for the functions $A$, $B$, $C$, and $V$.
Now we reconsider the metric given in Eq. (15) for a static, spherically symmetric stellar object harboring a wormhole, which reads
$ds^{2}=$ $-e^{2\nu(r)}dt^{2}+e^{2\lambda(r)}dr^{2}$ (22)
$+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\varphi^{2}\right).$
We employ an ansatz
$e^{2\lambda(r)}=\left(1-\frac{r_{0}}{r}\right)^{-1},$ (23)
where $r_{0}$ denotes the radius of the wormhole throat. Furthermore, for $r\geq R_{s}$, where $R_{s}$ is the surface radius, the exterior solution satisfies
$e^{2\nu(r)}=e^{-2\lambda(r)}=1-\frac{r_{0}}{r}.$ (24)
The ansatzes shown by Eqs. (23) and (24) have also been used in the ad hoc formalism within Ref. Nojiri2024 .
Note that even for the magnetized configuration of NSs, Eq. (24) can still be used as a good approximation, since the pressure and energy density associated with the magnetic field in the exterior region of magnetars are negligible (for $B_{s}=10^{15}\,\mathrm{G}$, both the magnetic energy density and pressure are of the order of $10^{-5}\,\mathrm{MeV\,fm^{-3}}$). Therefore, the exterior metric remains a valid approximation. A similar treatment is also employed in Refs. Pattersons2024a ; Mallick2014 .
For the WH+NS system, near the throat $r\sim r_{0}$, the metric function $e^{2\lambda(r)}$ behaves as Nojiri2024
$e^{2\lambda(r)}\sim\frac{r_{0}}{r-r_{0}}e^{2\lambda_{0}},$ (25)
where $\lambda_{0}$ is a constant. The radial coordinate $r$ can be redefined as Nojiri2024
$l=2\sqrt{r_{0}(r-r_{0})}.$ (26)
Near the throat, the metric becomes
$ds^{2}$ $\sim$ $-e^{2\nu(l)}dt^{2}+e^{2\lambda_{0}}dl^{2}$ (27)
$+r_{0}^{2}(d\theta^{2}+\sin^{2}\theta d\varphi^{2}).$
From the ansatz given in Eq. (23), one immediately obtains $\lambda_{0}=0$. According to Ref. Nojiri2024 , although the new radial coordinate $l$ is defined to be positive, the metric with the new radial coordinate $l$ admits an analytic continuation to the region $l<0$. We assume that the region $l>0$ corresponds to our spacetime, while the region $l<0$ represents another spacetime, which is smoothly connected to ours through the wormhole.
The EMT for anisotropic neutron fluid with chaotic magnetic field $\mathcal{T}_{\mu\nu}$ writes
$\mathcal{T}_{\mu\nu}\equiv T_{\text{matter}\>\mu\nu}+T_{\mathcal{B}\>\mu\nu},$
Note that the function $B$ appearing in Eq. (LABEL:action) can be determined by allowing for a time-dependent metric: in that case the $(t,r)$ component of the field equations does not vanish but yields $B=\frac{2\dot{\lambda}}{\kappa^{2}r}$. Since in the present time-independent metric $\dot{\lambda}=0$, we therefore obtain $B=0$.
From Eqs. (LABEL:tt)-(LABEL:thetatheta), we have
${\color[rgb]{0,0,0}\rho_{\text{tot}}=\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}+V+\varrho,}$ (36)
${\color[rgb]{0,0,0}p_{r,\text{tot}}=\frac{A}{2}e^{-2\nu}+\frac{C}{2}e^{-2\lambda}-V+P_{r},}$ (37)
${\color[rgb]{0,0,0}p_{t,\text{tot}}=\frac{A}{2}e^{-2\nu}-\frac{C}{2}e^{-2\lambda}-V+P_{t}.}$ (38)
After substituting the expressions of $A$, $C$, and $V$ from Eqs. (LABEL:A)-(LABEL:V) into Eqs. (36)-(38) we obtain
${\color[rgb]{0,0,0}\rho_{\text{tot}}=0},$ (39)
${\color[rgb]{0,0,0}p_{r,\text{tot}}=\frac{2r(r-r_{0})\nu^{\prime}-r_{0}}{\kappa^{2}r^{3}},}$ (40)
${\color[rgb]{0,0,0}p_{t,\text{tot}}=\frac{2r^{2}(r-r_{0})(\nu^{\prime\prime}+\nu^{\prime 2})+r(2r-r_{0})\nu^{\prime}+r_{0}}{2\kappa^{2}r^{3}}.}$ (41)
It is important to emphasize that the absence of explicit anisotropic or magnetic-field terms in the final expressions for the energy density and pressures does not imply that anisotropy or the magnetic field are physically absent from the system. Rather, within the present effective description, the total energy density, the total radial pressure, and the total tangential pressure admit the same functional form in terms of the metric functions, independently of whether anisotropy and/or magnetic fields are included in the matter sector.
As a consequence, the traversability conditions are governed by the total EMT, and remain unchanged both in the presence and in the absence of anisotropy and magnetic fields. In this sense, anisotropy and the magnetic field do not explicitly modify the traversability criteria, although they contribute implicitly to the effective geometry through the construction of the metric.
At the limit $r\rightarrow r_{0}$, we have
$(\rho_{\text{tot}}+p_{r,\text{tot}})$ $\;\longrightarrow\;$ $-\frac{1}{\kappa^{2}r_{0}^{\,2}},$
$(\rho_{\text{tot}}+p_{t,\text{tot}})$ $\;\longrightarrow\;$ $\frac{1+r_{0}\,\nu^{\prime}(r_{0})}{2\kappa^{2}r_{0}^{\,2}},$
which explicitly shows that the NEC is violated in the vicinity of the wormhole throat, and that our wormhole is traversable.
An important remark from Ref. Nojiri2024 should be emphasized: the violation of the energy conditions usually induces instabilities in the configuration. For instance, the sound speed may exceed the speed of light. In the wormhole spacetime constructed from the two-scalar model used in this work, however, such instabilities do not arise. This is because the two scalar fields are nondynamical; they neither propagate nor fluctuate. Consequently, no sound mode is generated from oscillations of the effective fluid produced by the scalar fields, even though this effective fluid violates the energy conditions.
4 Non-magnetized configuration of WH+NS systems
Before discussing the WH+NS configurations, we briefly clarify how the corresponding pure NS solutions are obtained. The NS configurations considered in this work are constructed using the standard TOV equations for a spherically symmetric relativistic fluid, without a wormhole throat. These solutions serve solely as reference configurations, allowing for a direct comparison between ordinary NSs and the WH+NS systems discussed below. No modification of the standard TOV formalism is assumed for the pure NS case.
From Eq. (LABEL:fieldscalar), for the WH+NS model considered here, one obtains
$0$ $=$ $\frac{1}{2}A_{\chi}e^{-2\nu}+\frac{1}{2}C_{\chi}e^{-2\lambda}$ (42)
$+Ce^{-2\lambda}\left(\frac{2}{r}+\nu^{\prime}-\lambda^{\prime}\right)-V_{\chi}.$
Recalling that $\chi=r$, Eq. (42) can be written as
$0$ $=$ $\frac{1}{2}\frac{dA}{dr}e^{-2\nu}+\frac{1}{2}\frac{dC}{dr}e^{-2\lambda}$ (43)
$+Ce^{-2\lambda}\left(\frac{2}{r}+\nu^{\prime}-\lambda^{\prime}\right)-\frac{dV}{dr}.$
By using the expression of $A$, $C$, and $V$ shown by Eqs. (LABEL:A)-(LABEL:V), we are left with
$\nu^{\prime}(P_{r}+\varrho)+P_{r}^{\prime}+\frac{2\sigma}{r}=0,$ (44)
which is the conservation law for anisotropic fluid.
It is worth emphasizing that the conservation equation (44) is not imposed as an independent assumption for the neutron fluid alone. Instead, it emerges naturally from the equations of motion obtained by varying the action with respect to the auxiliary scalar fields. This procedure guarantees the internal consistency of the constrained scalar-gravity system that underlies the WH+NS construction.
Physically, this approach ensures that the presence of auxiliary scalar fields, introduced to generate the effective exotic matter sector required by the wormhole geometry, does not spoil the standard conservation properties of the ordinary matter. Although the scalar fields contribute implicitly through the geometry and the effective EMT, the resulting conservation law involves only the ordinary matter and magnetic field variables. In this sense, the conservation equation should be interpreted as a consistency condition of the full system rather than as an externally imposed constraint.
For non-magnetized configuration, we have
$\nu^{\prime}(p_{r}+\rho)+p_{r}^{\prime}+\frac{2\sigma}{r}=0.$ (45)
By using the anisotropy profile shown by Eq. (7), Eq. (45) can be integrated to give
$\nu=-\left(2-\frac{1}{h}\right)\int\frac{dp_{r}}{\rho+p_{r}}.$ (46)
Note that in this work we use the polytropic equation of state (EOS), which reads Dzhunushaliev2012 ; Nojiri2024 ; daSilva2021 ; Godani2024 ; Ray2007
$p_{r}=K\rho^{1+\frac{1}{n}}.$ (47)
In this paper, we adopt the same parameters as those used in Ref. Nojiri2024 , namely $K=100~\mathrm{km}^{2}$ and $n=1$. Therefore, the EOS reduces to $p_{r}=K\rho^{2}$. For the matter-energy density profile, we also follow Ref. Nojiri2024 and employ a Tolman–VII-like profile, i.e.
$\rho=\rho_{c}\left[1-\left(\frac{r-r_{0}}{R_{s}-r_{0}}\right)^{2}\right],$ (48)
where $\rho_{c}$ denotes the energy density of the NS matter at $r=r_{0}$.
With straightforward integration of Eq. (45), one finds
$\nu=\nu_{c}-\left(4-\frac{2}{h}\right)\ln{\left(1+K\rho\right)},$ (49)
where $\nu_{c}$ is the integration constant. Consequently, the temporal metric function reads
$e^{2\nu}=\frac{e^{2\nu_{c}}}{(1+K\rho)^{8-\frac{4}{h}}}.$ (50)
At $r=R_{s}$, the energy density vanishes, $\rho(R_{s})=0$, and the temporal metric function satisfies $e^{2\nu(R_{s})}=1-\dfrac{r_{0}}{R_{s}}$. Substituting this condition into Eq. (50) yields
$e^{2\nu_{c}}=1-\frac{r_{0}}{R_{s}}.$ (51)
We can also compute the radial derivative of Eq. (50) at $r=R_{s}$. Using $\rho(R_{s})=0$ and $\rho^{\prime}(R_{s})=-2\rho_{c}/(R_{s}-r_{0})$ for the Tolman–VII-like profile, one obtains.
$(e^{2\nu})^{\prime}|_{r=R_{s}}=\left(16-\frac{8}{h}\right)\frac{K\rho_{c}}{R_{s}}.$ (52)
As $(e^{2\nu})^{\prime}|_{r=R_{s}}=\dfrac{r_{0}}{R_{s}^{2}}$, we obtain the relation
$r_{0}=\left(16-\frac{8}{h}\right)K\rho_{c}R_{s}.$ (53)
The metric given by Eq. (22) implies that the ADM mass $M$ satisfies
$M=\frac{r_{0}}{2}=\left(8-\frac{4}{h}\right)K\rho_{c}R_{s}.$ (54)
Generally, the total mass of the WH+NS systems is given by Nojiri2024
$M(r)=\frac{r_{0}}{2}+\int_{r_{0}}^{r}\rho_{\text{tot}}(r^{\prime})r^{\prime 2}dr^{\prime}.$ (55)
Since $\rho_{\text{tot}}=0$, the total mass is only taken from the ADM mass contribution, i.e. Eq. (54). It is also important to highlight that for the isotropic case where $h=1$, the calculations recover to the one presented in Ref. Nojiri2024 .
(a) Approach 1
(b) Approach 2
Figure 1: Mass–radius relation of the pure NSs and the WH+NS systems for different values of $h$. Panels (a) and (b) correspond to Approach 1 and Approach 2, respectively.
(a) Approach 1
(b) Approach 2
Figure 2: Radii of the stellar surface and the wormhole throat in the WH+NS configurations as functions of the ADM mass for different values of $h$. Panels (a) and (b) correspond to Approach 1 and Approach 2, respectively.
(a) Approach 1
(b) Approach 2
Figure 3: Compactness of the WH+NS systems as a function of the ADM mass for different values of $h$. Panels (a) and (b) correspond to Approach 1 and Approach 2, respectively.
Fig. 1 presents the mass-radius (MR) relations of the WH+NS configurations for several values of the parameter $h$. For comparison, we also display the MR curve of pure neutron stars (NSs), allowing the impact of the wormhole component to be directly identified.
To facilitate a consistent comparison with pure NS configurations, we adopt the following two approaches. Approach 1: for a given central energy density $\rho_{c}$, the radius of the surface, $R_{s}$, is assumed to coincide with the radius of an ordinary NS obtained at the same $\rho_{c}$. Approach 2: for a given central energy density $\rho_{c}$, the total mass of the system, $M$, is taken to be equal to that of an ordinary NS computed at the same $\rho_{c}$.
Both approaches have also been employed in Ref. Nojiri2024 , which serves as a methodological predecessor to the procedure adopted in this work. In the present analysis, we consider $h=1,;1.5,;2$.
For all considered values of $h$ within Approach 1, as we can see in Fig. 1(a), the mass of the WH+NS systems can exceed that of the secondary compact object in GW190814, indicating that the resulting configurations are exceptionally massive. Similar mass enhancement has already been reported for isotropic systems in Ref. Nojiri2024 . However, our inclusion of anisotropy allows the formation of even more massive configurations, accompanied by a larger surface radius.
Fig. 1(b) presents the MR relations obtained using Approach 2. Since pure polytropic NSs do not reach the large masses found in the configurations constructed under Approach 1, the WH+NS systems obtained via Approach 2 naturally cannot achieve comparable masses. As a consequence, the radii of the WH+NS configurations become significantly smaller and may even fall below $5\,\mathrm{km}$. We note that such objects, which appear externally similar to ordinary NSs but possess very small radii, could constitute a distinctive potential observational imprint should such configurations be realized in nature.
Fig. 2(a) shows the comparison between the radius of the WH+NS system and the corresponding wormhole throat radius obtained using Approach 1. We emphasize that the coincidence of the wormhole throat radii observed in Figs. 2(a) and (b) is exact. This is a direct consequence of the geometric relation $r_{0}=2M$, cf. Eq. (44). Since the ADM mass $M$ is used as the horizontal axis in the figures, the functional dependence $r_{0}(M)$ is fixed and identical for all configurations.
The role of the anisotropy parameter $h$ and other matter parameters is not to modify the relation between $r_{0}$ and $M$, but to determine the values of the ADM mass that are realized by physically admissible solutions. Different values of $h$ therefore lead to different mass ranges and, consequently, to different accessible throat radii through the relation $M=r_{0}/2$, while preserving the exact coincidence of the $r_{0}(M)$ curves.
The wormhole throat radii obtained using Approach 2, shown in Fig. 2(b), exhibit a trend similar to that found in Approach 1. The primary difference between the two approaches appears in the overall mass and radius scales: Approach 1 yields systematically larger masses and radii, consistent with the corresponding MR relations obtained under that approach.
As mentioned earlier, the WH+NS system can reach masses that may even exceed the mass of the secondary compact object of GW190814. This fact implies that the system can become very massive and, consequently, highly compact, as illustrated in Fig. 3. The compactness $\mathcal{C}=M/R_{s}$ can even enter the region typically associated with ultracompact objects. The range of $\mathcal{C}$ for ultracompact object is $0.33<\mathcal{C}<0.44$ ByonBurikham ; ByonAnto . Recently, research on ultracompact objects has attracted significant interest (see, e.g., Refs. ByonBurikham ; ByonAnto ; Prasetyo2021 ; Prasetyo2021b ; Fauzi2025 ; Carballo2022 ; Benetti2025 ; Urbano2019 ; Prasetyo2023 ; Hidalgo2021 ; Tasinato2022 ; Kleihaus2020 ; Zhu2020 ). Those objects possess radii smaller than $3M$, i.e., inside the photon-sphere radius. In such regimes, gravitational-wave echoes may occur, and the corresponding echo time can be computed Sakti2021 , thereby providing a potential observational imprint of such systems. The fact that the system considered in this work can enter the ultracompact region is therefore a notable result, as this regime cannot be achieved by ordinary polytropic NSs.
(a) Approach 1
(b) Approach 2
Figure 4: Surface redshift of the WH+NS systems as a function of the ADM mass for different values of $h$. Panels (a) and (b) correspond to Approach 1 and Approach 2, respectively.
(a) Approach 1
(b) Approach 2
Figure 5: Echo time of the WH+NS systems as a function of the ADM mass for different values of $h$. Panels (a) and (b) correspond to Approach 1 and Approach 2, respectively.
Some physical quantities can serve as potential observational imprints of WH+NS configurations, one of which is the surface redshift $z$. Mathematically, it is defined as Sakti2021
$z=\frac{1}{e^{\nu(R_{s})}}-1.$ (56)
For ordinary NSs with typical masses $M=1.2\text{--}2.0\,M_{\odot}$ and radii $R_{s}=10\text{--}14$ km, the expected surface redshift falls within the range $z\simeq 0.1\text{--}0.5$. Cottam et al. Cottam2002 reported a gravitational redshift of $z=0.35$ from 28 bursts of the low-mass X-ray binary EXO 0748-676, consistent with theoretical expectations for regular NSs.
Fig. 4 shows the surface redshift of the non-magnetized WH+NS systems as a function of the ADM mass for various values of $h$. The results indicate that the redshift can exceed $z=1$, meaning that the predicted $z$ lies well above typical values for ordinary NSs. This suggests that the configurations considered here behave as exotic compact objects. We also find that both Approach 1 and Approach 2 yield comparable values of the parameter $z$. However, these values appear in different mass ranges, reflecting the fact that the two approaches lead to systematically different mass distributions. This indicates that the parameter $z$ is relatively insensitive to the choice of approach, while the associated mass scale depends strongly on the adopted construction scheme. Furthermore, in both cases, increasing anisotropy tends to reduce the resulting value of $z$. Hence, the disappearance of anisotropy with positive $h$ values prolongs the wavelength that is measured by the observer at infinity. This finding serves as a promising indicator for the possible detection of such WH+NS systems.
Another promising potential observational imprint of the WH+NS systems is the echo time. By assuming that the wormhole is connected to an identical object residing in another spacetime—following the setup considered in Ref. Lai2025 —gravitational waves can propagate through the wormhole mouths and become trapped between the two photon spheres. This process gives rise to gravitational echoes (see Fig. 1 of Ref. Lai2025 for a schematic illustration).
A distinctive feature of such systems is that the resulting signal is expected to have an extremely short duration and to lack a clearly identifiable inspiral phase, which is commonly observed in standard compact binary mergers. In this scenario, the detected gravitational-wave signal essentially consists of the echo component alone, originating from waves repeatedly traversing the wormhole throat.
Within this framework, the echo time $\tau$ can be expressed as Lai2025
$\tau=2\int_{r_{0}}^{R_{p}}e^{\lambda-\nu}dr.$ (57)
Here, $R_{p}$ is the radius of the photon sphere. At the photon sphere, the photon must move at constant values of $r$ and $\theta$. This condition gives
$\left(\frac{d\varphi}{dt}\right)^{2}=\frac{1-\frac{r_{0}}{r}}{r^{2}\sin^{2}\theta}.$ (58)
On the other hand, from geodesic equation, one can obtain
$\left(\frac{d\varphi}{dt}\right)^{2}=\frac{r_{0}}{2r^{3}\sin^{2}\theta}.$ (59)
By combining Eqs. (58) and (59), one obtains $r=3r_{0}/2=3M$, which corresponds to the location of the photon sphere. It is important to emphasize that this result is model dependent. For instance, in alternative compact object configurations such as dark energy stars with phantom fields, the photon sphere radius $R_{p}$ can deviate from the value $3M$. In such scenarios, the upper limit of the integral in Eq. (57) is consequently modified, as demonstrated in Ref. Sakti2021 .
Since the functional form of $e^{2\nu(r)}$ differs between the regions $r_{0}\leq r<R_{s}$ and $r\geq R_{s}$, the integration in Eq. (57) must be performed piecewise. Accordingly, we split the integration domain into two intervals: from $r_{0}$ to $R_{s}$, and from $R_{s}$ to $R_{p}$.
In this work, the integral over the inner region, $r_{0}\leq r<R_{s}$, is evaluated numerically due to the complicated form of $e^{2\nu(r)}$. On the other hand, the integral over the outer region, $R_{s}\leq r\leq R_{p}$, admits a closed-form analytical expression, given by
$\int_{R_{s}}^{R_{p}}e^{\lambda-\nu}dr=(R_{p}-R_{s})+r_{0}\ln\left(\frac{R_{p}-r_{0}}{R_{s}-r_{0}}\right).$ (60)
Fig. 5 displays the resulting echo time as a function of the system’s total mass. The echo time is found to lie within a very short interval, namely of order $10^{-1}\,\mathrm{ms}$ for Approach 1 and between $10^{-2}$ and $10^{-1}\,\mathrm{ms}$ for Approach 2. This behavior is expected: the presence of a wormhole at the center effectively creates a “hollow” region in the geometry, such that the echo signal is abruptly transported to a different branch of the spacetime. As a consequence, the optical path length decreases. This effect becomes even more pronounced when the wormhole throat radius approaches the stellar surface radius.
From Fig. 5, it can also be seen that in Approach 1, configurations with the same mass but larger values of $h$ exhibit shorter echo times. In Approach 2, although configurations with identical masses for different $h$ cannot be directly compared on the same plot, one may imagine extending each curve following its visible trend; under this qualitative continuation, larger values of $h$ would again correspond to shorter echo times. This situation–at least for Approach 1–is consistent with the behavior of the surface redshift in Fig. 4, where smaller $h$ at fixed mass yields a larger redshift $z$. Physically, a larger $z$ indicates stronger gravitational trapping, making it more difficult not only for light to escape but also for the echo signal to climb out of the gravitational well.
5 Magnetized configuration of WH+NS systems
Before discussing the WH+NS configurations, we clarify the role of the parameter $\rho_{0}$ used throughout this work. Here, $\rho_{0}$ denotes the central energy density of the corresponding non-magnetized pure NS configurations and is treated as a scanning parameter.
The range $10^{-4}\leq\rho_{0}\leq 10^{4}\,\mathrm{MeV\,fm^{-3}}$ is chosen to encompass the physically relevant domain of central energy densities typically considered in NS structure calculations, from low-density solutions to very dense cores. Thus, $\rho_{0}$ must exist within this range.
For each value of the anisotropy parameter $h$, the value of $\rho_{0}$ is determined numerically from the corresponding pure neutron-star solutions. Specifically, we solve the standard TOV equations for non-magnetized neutron stars by varying the central energy density $\rho_{c}$ and identify the value of $\rho_{c}$ that yields the maximum-mass configuration.
For the specific values of the anisotropy parameter considered in this work, the maximum-mass pure NS configurations are found at $\rho_{0}=3240.46$ MeV fm-3 for $h=1$, $\rho_{0}=2008.27$ MeV fm-3 for $h=1.5$, and $\rho_{0}=1459.81$ MeV fm-3 for $h=2$, respectively. All these values lie well within the range $10^{-4}\leq\rho_{0}\leq 10^{4}\,\mathrm{MeV\,fm^{-3}}$ adopted in our analysis.
The EMT for magnetized configuration is shown by Eq. (LABEL:EMTmagnetized). The conservation law gives
$\nu=-\left(2-\frac{1}{h}\right)\int\frac{dP_{r}}{\varrho+P_{r}}.$ (61)
With a little calculation, we obtain
Considering that we face two possibilities for the second term on the right-hand side of Eq. (LABEL:integrallongI1), we decide to perform the calculation using four ranges of surface magnetic field strengths, while keeping the range of $\rho_{0}$ fixed between $1\times 10^{-4}$ and $1\times 10^{4}$ MeV fm-3. The four ranges are summarized in Table 1.
Table 1: Four considered ranges of the surface magnetic field $B_{s}$ and the corresponding core magnetic field $B_{0}$ considered in this work.
Fig. 6 $B_{s}$ (G) $B_{0}$ (G)
(a) $10^{12}-10^{13}$ $3\times 10^{15}$
(b) $10^{13}-10^{14}$ $3\times 10^{16}$
(c) $10^{14}-10^{15}$ $3\times 10^{17}$
(d) $10^{15}-10^{16}$ $3\times 10^{18}$
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 6: Three-dimensional plot of the discriminant $D=\beta^{2}-4\alpha\gamma$ as a function of $\rho_{0}$ and $B_{s}$ for (a) $B_{0}=3\times 10^{15}$ G, (b) $B_{0}=3\times 10^{16}$ G, (c) $B_{0}=3\times 10^{17}$ G, and (d) $B_{0}=3\times 10^{18}$ G.
The choice of these values is motivated by the fact that the typical surface magnetic field of NSs lies in the range $10^{12}$$10^{15}$ G. Furthermore, according to the simulations reported in Ref. Lopes2020 , a surface field of $B_{s}=10^{15}$ G corresponds to a core magnetic field of $B_{0}=3\times 10^{18}$ G. Therefore, in our analysis we restrict the order of magnitude of $B_{0}$ to be at most three orders of magnitude larger than $B_{s}$. For the values of $\rho_{0}$ used in this work, our non-magnetized pure NS calculations give $\rho_{0}=3240.46~\text{MeV fm}^{-3}$ for $h=1$, $\rho_{0}=2008.27~\text{MeV fm}^{-3}$ for $h=1.5$, and $\rho_{0}=1459.81~\text{MeV fm}^{-3}$ for $h=2$. We see that all these values remain inside the range $\times 10^{-4}1×1$$\times 10^{4}~\text{MeV fm}^{-3}$. Thus, all values of $\rho_{0}$, $B_{s}$, and $B_{0}$ used throughout this study lie within the parameter ranges described above. The small value of $\rho_{0}$ is intentionally taken so that the chosen second term of the right-hand side of Eq. (LABEL:integrallongI1) remains applicable even in the small-order regime, without requiring concern over how small $\rho_{0}$ is.
(a) Approach 1
(b) Approach 2
(c) Approach 1
(d) Approach 2
Figure 7: MR relations of the pure NSs and the WH+NS systems for different values of the anisotropy parameter $h$. Panels (a) correspond to $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, (b) $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2, (c) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, and (d) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2.
(a) Approach 1
(b) Approach 2
(c) Approach 1
(d) Approach 2
Figure 8: Radii of the stellar surface and the wormhole throat in the WH+NS configurations with $B_{s}=10^{15}$ G and $B_{0}=10^{18}$ G as functions of the ADM mass for different values of $h$. Panels (a) correspond to $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, (b) $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2, (c) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, and (d) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2.
(a) Approach 1
(b) Approach 2
(c) Approach 1
(d) Approach 2
Figure 9: Compactness of the WH+NS configurations with $B_{s}=10^{15}$ G and $B_{0}=10^{18}$ G as functions of the ADM mass for different values of $h$. Panels (a) correspond to $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, (b) $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2, (c) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, and (d) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2.
(a) Approach 1
(b) Approach 2
(c) Approach 1
(d) Approach 2
Figure 10: Surface redshift of the WH+NS configurations with $B_{s}=10^{15}$ G and $B_{0}=10^{18}$ G for different values of $h$. Panels (a) correspond to $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, (b) $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2, (c) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, and (d) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2.
(a) Approach 1
(b) Approach 2
(c) Approach 1
(d) Approach 2
Figure 11: Echo time of the WH+NS configurations with $B_{s}=10^{15}$ G and $B_{0}=10^{18}$ G as functions of the ADM mass for different values of $h$. Panels (a) correspond to $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, (b) $B_{s}=10^{15}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2, (c) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 1, and (d) $B_{s}=10^{16}\,\mathrm{G}$ and $B_{0}=10^{18}\,\mathrm{G}$ with Approach 2.
In Fig. 6, the gray region corresponds to the range $0\leq D<10^{-40}$, while the black region represents negative values of $D$. Note that $1\times 10^{-40}$ is the maximum threshold of numerical precision in our computation. We observe that no gray or black regions appear in the figure, indicating that the discriminant remains positive and always above $1\times 10^{-40}$ throughout all computed regimes. This shows that we can use the logarithmic function for the second term on the right-hand side of Eq. (LABEL:integrallongI1), instead of the inverse tangent function. Now Eq. (LABEL:integrallongI1) reads
In Appendix B, we show that for the isotropic configuration with the limit $\mathcal{B}\rightarrow 0$, $\left(e^{2\nu}\right)^{\prime}$ is reduces to the expression of $\left(e^{2\nu}\right)^{\prime}$ shown by Eq. (24) in Ref. Nojiri2024 .
At the surface, $\left(e^{2\nu}\right)^{\prime}\big|_{r=R_{s}}=\frac{r_{0}}{R_{s}}$, so we have
$r_{0}=\left[\frac{4K\left(2-\frac{1}{h}\right)\rho_{c}}{\alpha\gamma}(\beta-1)+\frac{\rho_{c}zB_{s}}{3\pi\gamma}\right]R_{s}.$ (76)
Furthermore, by substituting the expressions of $\alpha$, $\beta$, $\gamma$, and $z$ in Eq. (LABEL:alphabetagamma) into Eq. (76), we are left with
$r_{0}=2\left[\frac{4K\left(2-\frac{1}{h}\right)}{K+\frac{B_{0}^{2}}{6\pi\rho_{0}^{2}}}+1\right]\frac{\rho_{c}B_{0}}{\rho_{0}B_{s}}R_{s}.$ (77)
From Eq. (77), we can easily obtain the ADM mass
$M=\left[\frac{4K\left(2-\frac{1}{h}\right)}{K+\frac{B_{0}^{2}}{6\pi\rho_{0}^{2}}}+1\right]\frac{\rho_{c}B_{0}}{\rho_{0}B_{s}}R_{s}.$ (78)
A remark is stressed to be emphasized in order concerning the limit of vanishing magnetic field. In the magnetized configurations, the magnetic field modifies the geometric structure of the solution. As a result, the limit $\mathcal{B}\to 0$ is non-uniform: setting $\mathcal{B}=0$ directly at the level of intermediate quantities does not, in general, reproduce the non-magnetized configuration obtained by solving the full system.
As shown explicitly in A and B, the metric function $e^{2\nu}$ and its radial derivative $\left(e^{2\nu}\right)^{\prime}$ smoothly reduce to their non-magnetized counterparts when the limit $\mathcal{B}\to 0$ is taken at the level of the full metric functions. This point is crucial because the wormhole throat radius $r_{0}$ is determined from $\left(e^{2\nu}\right)^{\prime}$ evaluated at the surface $r=R_{s}$, where the energy density vanishes. At this stage, $\left(e^{2\nu}\right)^{\prime}$ contains the coefficients $\alpha$, $\beta$, and $\gamma$, which depend on the magnetic field and appear both in the numerator and the denominator, leading to the finite expression given in Eq. (77).
If the limit $\mathcal{B}\to 0$ is taken prematurely at the level of individual quantities, unphysical results may arise. For instance, the coefficient $e^{2\nu_{c}}$ appearing in the magnetized solution, shown by Eq. (LABEL:integrationconstantmagnetized), explicitly depends on the magnetic-field parameter $\gamma$ and would vanish identically if the limit were applied naively, causing the metric function $e^{2\nu}$ itself to collapse. This pathology is avoided once the limit is taken at the level of the full metric function, where the cancellation of magnetic-field–dependent terms occurs consistently, and the system is reduced to non-magnetized configuration.
This behavior shows that the magnetic field genuinely modifies the geometric structure of the system and that the non-magnetized configuration is recovered only when the limit is applied at the appropriate structural level. The magnetized and non-magnetized solutions therefore belong to distinct branches of the solution space, which justifies treating them separately in this work.
Note that in this work, regarding the influence of the magnetic field on the EOS, we follow the approach used in Refs. Mallick2014 ; Konno1999 ; Konno2000 ; Rizaldy2018b ; Rizaldy2019 , where the magnetic field contribution to the EOS is taken to be negligible. In particular, Refs. Konno1999 ; Konno2000 also employed a polytropic EOS in their analysis. The authors of Ref. Mallick2014 argued that the dominant magnetic effects arise from the additional stress and pressure terms, while the contribution from magnetization remains subdominant even for extremely strong magnetic fields, especially when such high central fields tend to drive the star toward instability. They used magnetic fields of the order of $10^{18}$ G in their analysis. The neglect of Landau quantization effects on the EOS has also been discussed in Ref. Sinha2013 .
In the numerical calculations of magnetized WH+NS configurations we consider two representative combinations of magnetic fields: (i) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=10^{15}\,\mathrm{G}$, and (ii) $B_{0}=3\times 10^{18}\,\mathrm{G}$ and $B_{s}=10^{16}\,\mathrm{G}$. Although the typical surface magnetic field of NSs lies in the range $10^{12}$$10^{15}\,\mathrm{G}$ as mentioned in Sec. 1, Ref. Gomes2019 reports that the limiting magnetic moment of NSs corresponds to a surface field of order $B_{s}\sim 10^{16}\,\mathrm{G}$. For this reason we also include $B_{s}=10^{16}\,\mathrm{G}$ among the cases examined in our numerical calculations. Configurations (i) and (ii) lie within the parameter ranges shown in Fig. 6. The anisotropy parameter is chosen to be the same as in the non-magnetized case, namely h=1, 1.5,h=1,\ 1.5, and $2$.
The MR relations obtained using Approach 1 are shown in Fig. 7(a), while those corresponding to Approach 2 are presented in Fig. 7(b). The two approaches lead to markedly different mass and radius ranges. Approach 1 yields configurations with large masses and radii, whereas Approach 2 produces systems with extremely small masses and radii. For this reason, in the case of Approach 2, a direct comparison with ordinary NSs is not meaningful, as the mass–radius ranges differ by several orders of magnitude. Including the ordinary NS curves would render the WH+NS MR relations visually indistinguishable in the plot.
Another important point is that observational constraints indicate that the typical radius of neutron stars lies in the range of approximately 1010-$11.5$ km ozel2016 , while the canonical neutron-star mass is $1.4\,M_{\odot}$ Pattersons2025 . Therefore, the very large radii in the range 20209090 km obtained in Fig. 7 are incompatible with current observational evidence, at least at the present stage.
However, it should be emphasized that these large-radius configurations arise from the use of a polytropic EOS in the low density range, for which such extended radii are mathematically allowed. The absence of observational counterparts for these configurations is further supported by the fact that they correspond to extremely small ADM masses, approaching zero and thus lying far below the canonical mass of $1.4\,M_{\odot}$.
In contrast, within Approach 1, WH+NS systems with similarly large radii can be realized only at the cost of extremely large masses. In particular, for $h=2$, the total mass obtained in Approach 1 can exceed $8\,M_{\odot}$, significantly larger than the mass of the secondary compact object in GW190814. Moreover, increasing the surface magnetic field further enhances both the mass and radius of the system: for $B_{s}=10^{16}\,\mathrm{G}$, the resulting configurations can attain total masses exceeding $14\>M_{\odot}$. This highlights the extreme and observationally challenging nature of such configurations.
Approach 2 yields configurations with radii between 0-$0.7$ km, as shown in Fig. 7(b) and (d). The corresponding MR curves exhibit significantly steeper trends compared to those obtained with Approach 1. For $B_{s}=10^{16}\,\mathrm{G}$, systems with low anisotropy parameter $h$ display the usual steep MR behavior, with only a mild flattening forming a shallow hump. In contrast, for anisotropic cases, a pronounced hump gradually develops as $h$ increases, while the steep branch becomes progressively shorter. In particular, for $h=2$, the maximum mass of the system no longer occurs along the steep branch of the MR curve, but instead is reached at the peak of the hump structure, although this peak remains relatively moderate and does not form an extreme maximum.
The behavior observed in Fig. 7 can be understood by recalling how the reference scales are fixed in the two approaches. In both cases, the calibration is performed with respect to the corresponding pure NS solutions. In Approach 1, the surface radius of the WH+NS system is fixed to coincide with the radius of the pure NS, while in Approach 2 the total ADM mass is fixed to the mass of the pure NS at the same central density.
In the non-magnetized configuration, the masses and radii of pure NSs lie within a moderate and physically familiar range. The obtain masses and radii of the WH+NS systems also lie within moderate range, and comparable to MR relation of the pure NS, as illustrated in Fig. 1.
The inclusion of a magnetic field qualitatively alters this picture. In magnetized configurations, the wormhole throat radius $r_{0}$, and therefore the ADM mass $M=r_{0}/2$, explicitly depend on the magnetic field parameters. At large radii, where the mass of a pure magnetized NS becomes very small, a WH+NS configuration constructed using Approach 1 necessarily develops an extremely large mass as the consequence of the Eq. (77). This leads to the appearance of very massive WH+NS systems, as shown in Figs. 7(a) and (c).
On the other hand, when using Approach 2, the ADM mass of the WH+NS system is constrained to remain equal to that of the pure magnetized NS. Under this condition, the relation $M=r_{0}/2$ forces the throat radius and the surface radius to shrink to very small values. The extremely small masses and radii observed in Figs. 7(b) and (d) therefore reflect the non-trivial interplay between the magnetic field and the imposed mass constraint, rather than any numerical instability.
Overall, the magnetic field gives rise to WH+NS configurations with MR properties that differ qualitatively from those of the non-magnetized case: Approach 1 naturally leads to extremely massive systems, whereas Approach 2 produces configurations with extremely small radii.
Fig. 8 shows the wormhole throat radius $r_{0}$ and surface radius $R_{s}$ as the functions of the ADM mass $M$ for different values of the anisotropy parameter $h$ and the different surface magnetic field $B_{s}$, within two construction approaches considered in this work. In panels (a) and (c), corresponding to Approach 1, we can obvioulsy see that the throat-radius curves coincide. In panel (b), corresponding to Approach 2, the solutions exist only within distinct mass intervals. As a result, the throat radius curves for $h=1.5$ and $h=2$ appear separated. This behavior does not indicate a modification of the functional relation $r_{0}(M)$, but rather reflects the restricted mass ranges over which the corresponding solutions are physically admissible. A different behavior is seen in panel (d), which also corresponds to Approach 2. Here, we can see that the curves of the throat radius are separated to several disconnected mass intervals. Within each interval the relation $r_{0}=2M$ is still satisfied, but the allowed ranges of the throat radius differ, which explains the clearly separated curves shown in the figure. It is important to note that anisotropy does not modify the geometric relation between $r_{0}$ and $M$, but it does affect the domain of existence of WH+NS solutions, with a particularly strong impact in Approach 2.
Fig. 9 shows the compactness of the WH+NS systems for the magnetized configurations considered in this work. We find that all curves can enter the ultracompact regime, indicating the existence of parameter ranges in which the echo time can be meaningfully evaluated. Systems with a higher surface magnetic field generally populate higher mass ranges; consequently, the ultracompact region is reached at larger masses compared to configurations with lower surface magnetic fields. Notably, the curves corresponding to $B_{s}=1\times 10^{16}\,\mathrm{G}$ within Approach 2 exhibit trends that are markedly different from the other cases. This behavior can be understood as a direct consequence of the distinct MR relations obtained in this configuration.
Another important point to emphasize is that, in the magnetized configurations, as mentioned before, Approach 2 yields systems with extremely small masses, even though their compactness values are comparable to those obtained using Approach 1. Objects with such properties are expected to be exceedingly difficult to detect observationally. Thus, Approach 2 is not relevant for obtaining physically detectable objects (at least at the present time). Nevertheless, they remain of theoretical interest, and for this reason we still include them in our calculations of the surface redshift $z$ and the echo time $\tau$.
The mathematical expression used to calculate the surface redshift is the same as that given in Eq. (56). From Fig. 10, we observe that the overall trend of the surface redshift closely resembles that found in the non-magnetized configurations, where larger values of the anisotropy parameter $h$ lead to smaller values of the surface redshift $z$. An exception occurs for the configuration with $B_{s}=1\times 10^{16}\,\mathrm{G}$ within Approach 2, in which each curve spans a different mass range. In this case, a direct point-by-point comparison between the curves is not meaningful; nevertheless, it can still be seen that smaller values of $h$ correspond to narrower mass ranges.
Although the qualitative behavior of $z$ remains similar to that of the non-magnetized case, it is important to note that the numerical range of the surface redshift values is also comparable to those obtained without magnetic fields. As expected, Approach 1 produces a much broader and higher mass range than the non-magnetized configurations, whereas Approach 2 yields a significantly smaller mass range in comparison.
Fig. 11 presents the echo time for each magnetized configuration obtained using Approach 1 and Approach 2. For Approach 1, both configurations with $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$ yield echo times of order $10^{-1}$ ms, which is the same order of magnitude as in the non-magnetized case. However, the echo times in the magnetized configurations are systematically larger than those obtained without magnetic fields. In the non-magnetized configuration, the longest echo time considered corresponds to $h=2$ and remains below $0.4$ ms. By contrast, for $B_{s}=1\times 10^{15}\,\mathrm{G}$ the maximum echo time at $h=2$ exceeds $0.6$ ms, while for $B_{s}=1\times 10^{16}\,\mathrm{G}$ it can exceed $0.9$ ms. This behavior can be attributed to the larger surface radii induced by the magnetic field, which increase the characteristic propagation distance of the echo signals.
On the other hand, Approach 2, shown in Figs. 11(b) and (d), yields extremely short echo times, of order $10^{-2}\,\mu\mathrm{s}$ for $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $10^{-1}\,\mu\mathrm{s}$ for $B_{s}=1\times 10^{16}\,\mathrm{G}$. This behavior can be understood as a consequence of the very small separation between the wormhole throat radius $r_{0}$ and the stellar surface $R_{s}$, which confines the echo propagation to a very short radial distance. We note that such short echo times would be extremely challenging to detect. Nevertheless, as emphasized earlier, Approach 2 is primarily of theoretical interest and is included to illustrate the range of possible behaviors allowed by the model.
In general, a direct comparison between the systems with $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$ is not straightforward, since the mass ranges over which echo signals exist differ between the two cases. An exception occurs for the $h=1$ configuration in Approach 1, where the admissible mass intervals overlap in the range $M\simeq 3.90$$4.17\,M_{\odot}$, allowing a meaningful comparison. The comparison is shown by Table 2.
Table 2: Echo time for magnetized WH+NS systems with $h=1$ for different masses. The comparison is shown for surface magnetic fields $B_{s}=1\times 10^{15}\,\mathrm{G}$ and $B_{s}=1\times 10^{16}\,\mathrm{G}$.
$M\,(M_{\odot})$ $\tau$ (ms)
$B_{s}=1\times 10^{15}\,\mathrm{G}$ $B_{s}=1\times 10^{16}\,\mathrm{G}$
3.90 0.24 0.20
3.97 0.25 0.21
4.03 0.26 0.22
4.12 0.27 0.23
4.17 0.29 0.24
Since we obtain the echo times in magnetized configurations that are larger than those in the non-magnetized case, one might expect the echo time to increase monotonically with the magnetic-field strength. However, Table 2 shows that, within the mass range where a direct comparison is possible, the systems with $B_{s}=1\times 10^{16}\,\mathrm{G}$ exhibit shorter echo times than those with $B_{s}=1\times 10^{15}\,\mathrm{G}$. In order to clarify the origin of such nontrivial behavior, we further analyze the dependence of the echo time on the magnetic field strength $\mathcal{B}$. To address this, we consider a toy model in which the magnetic field is taken to be constant from $r_{0}$ to $R_{s}$, and the pressure is assumed to be isotropic, such that the conservation law yields
$\nu$ $=$ $-2\int\frac{\rho d\rho}{\rho+p_{r}+\frac{\mathcal{B}^{2}}{6\pi}}$
$\nu$ $=$ $\nu_{c}-\ln\bigg[\rho(K\rho+1)+\frac{\mathcal{B}^{2}}{6\pi}\bigg]+\frac{1}{\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}$ (79)
$\times\ln\left[\frac{2K\rho+1-\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}{2K\rho+1+\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}\right].$
The entity $e^{2\nu}$ will be
$e^{2\nu}$ $=$ $e^{2\nu_{c}}\left(\frac{2K\rho+1-\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}{2K\rho+1+\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}\right)^{\frac{2}{\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}}$ (80)
$\times\left[\rho(K\rho+1)+\frac{\mathcal{B}^{2}}{6\pi}\right]^{-2}.$
By recalling $e^{2\nu(R_{s})}=1-\frac{r_{0}}{R_{s}}$, we can easily find
$e^{2\nu_{c}}$ $=$ $\left(\frac{1-\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}{1+\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}\right)^{-\frac{2}{\sqrt{1-4K\frac{\mathcal{B}^{2}}{6\pi}}}}$ (81)
$\times\left(1-\frac{r_{0}}{R_{s}}\right)\bigg(\frac{\mathcal{B}^{2}}{6\pi}\bigg)^{2}.$
Figure 12: Echo time as a function of $\mathcal{B}$.
With the $e^{2\nu}$ of the toy model in our hands, the echo time can be calculated using the same expression as in Eq. (57). In this stage of calculation, as in the previous cases, the integral over the inner region, $r_{0}\leq r<R_{s}$, is evaluated numerically. The integral over the outer region, $R_{s}\leq r\leq R_{p}$, on the other hand, also admits a closed-form analytical expression, given by the second term on the right-hand side of Eq. (60).
Table 3: Sets of input parameters used for calculating the echo time as a function of $\mathcal{B}$ in the toy model
Set $R_{s}$ (km) $r_{0}$ (km) $R_{p}$ (km) $\rho_{c}$ (MeV fm-3)
1 14.93 10.66 15.99 0.078
2 17.07 11.99 17.98 0.769
Here, we consider two sets of calculation inputs, as summarized in Table 3. The compactness of the system in Set 1 is $0.36$, while that of Set 2 is $0.35$. We emphasize that the mass and radii of the systems generally depend strongly on the magnetic field. However, since the present analysis is intended as a toy model, these dependencies are neglected. Our aim is solely to investigate the influence of the magnetic field on the echo time.
In Fig. 12, we observe that the echo time tends to remain nearly constant over a wide range of the magnetic field strength $\mathcal{B}$, taking values of the order of $10^{-1}$ ms. This magnitude is comparable to that obtained in the non-magnetized configuration. So it is clear that the echo time at $\mathcal{B}\rightarrow 0$ is consistent with the echo time shown by Fig. 5. The echo time exhibits a flat behavior up to $\mathcal{B}\sim 10^{16}$ G, beyond which it drops sharply.
Within this simplified toy-model analysis, the echo-time profile as a function of the magnetic field provides useful insight. In particular, it clearly indicates that the echo time may decrease for strong surface magnetic fields around $B_{s}=10^{16}$ G, offering a plausible explanation for the behavior we see in the magnetized configurations presented by Table 2.
6 Conclusion
In this study, we have constructed the formalism of anisotropic WH+NS systems, both in the absence and in the presence of magnetic fields. In general, the underlying model admits ghosts; however, these ghosts can be eliminated by imposing appropriate constraints through the introduction of Lagrange multiplier fields in the extended action. We find that the wormhole remains traversable irrespective of whether fluid anisotropy and/or magnetic fields are present, in the sense that the NEC is violated in the vicinity of the wormhole throat both with and without these effects.
We investigate the physical properties of the resulting configurations, including the ADM mass, stellar radius, wormhole throat radius, and compactness. In addition, we analyze potential observational imprints, namely the surface redshift and the gravitational-wave echo time. Throughout our analysis, we adopt two complementary approaches, following the methodology introduced in Ref. Nojiri2024 . In Approach 1, for a given central energy density $\rho_{c}$, the stellar radius is assumed to coincide with that of an ordinary NS obtained at the same $\rho_{c}$. In Approach 2, for the same $\rho_{c}$, the total mass of the WH+NS system is taken to be equal to that of an ordinary NS computed at the same central energy density.
We find that, within Approach 1, both non-magnetized and magnetized WH+NS configurations can become extremely massive. In particular, the total mass of the system can exceed that of the secondary compact object in GW190814, and may even surpass $8\,M_{\odot}$. In contrast, Approach 2 yields significantly lower masses by construction, since the total mass of the WH+NS system is fixed to be equal to that of an ordinary polytropic neutron star with the same central energy density. In the magnetized configurations of Approach 2, the resulting masses can even approach values close to zero.
Despite these differences in mass, the compactness of all configurations considered in this work can reach the ultracompact regime. Regarding the role of the magnetic field, we find that a larger surface magnetic field generally leads to larger masses and stellar radii. However, magnetized configurations obtained within Approach 2 are expected to be extremely difficult to observe, due to their very small masses and radii. Nevertheless, such configurations remain of theoretical interest.
In the context of the surface redshift, we find that all configurations considered in this work can exceed values of $z1.5$, which are significantly larger than those typically expected for ordinary neutron stars. Such large surface redshift values therefore provide a promising potential observational imprint of WH+NS systems. Furthermore, for a fixed configuration and at a given mass, larger values of the anisotropy parameter $h$ systematically lead to smaller surface redshift values.
The echo time for non-magnetized configurations is of the order of $10^{-2}$$10^{-1}$ ms. For magnetized configurations, the echo time in Approach 1 is of order $10^{-1}\,\mathrm{ms}$ and is larger than that of the corresponding non-magnetized configurations. In contrast, Approach 2 yields much shorter echo times for magnetized systems, of order $10^{-2}$$10^{-1}\,\mu\mathrm{s}$.
In addition, we construct a formulation of the echo time as an explicit function of the magnetic field, assuming a constant magnetic field throughout the system. The resulting behavior exhibits a flat behavior from $B\simeq 0$ up to $B\sim 10^{15}$ G, followed by a sharp decrease near $B\sim 10^{16}$ G. This trend explains why, within the mass range where a direct comparison is possible, the echo times associated with a surface magnetic field $B_{s}=1\times 10^{16}\,\mathrm{G}$ are smaller than those obtained for $B_{s}=1\times 10^{15}\,\mathrm{G}$.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, M. Lawrence Pattersons used ChatGPT (OpenAI) in order to support brainstorming and organization of some theoretical ideas of this work, and to improve the language and readability of the manuscript. All physical interpretations, model constructions, and final manuscript preparation were performed by the authors, who take full responsibility for the content. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the paper.
Acknowledgments
MLP gratefully acknowledges Byon Jayawiguna for precious discussions on the echo time. MLP is also financially supported by Indonesia Endowment Fund for Education Agency (LPDP). HLP was funded by the postdoctoral program in BRIN (National Research and Innovation Agency), Indonesia. MFARS is supported by the Second Century Fund (C2F) and C2F research abroad scholarship. We would like to thank the members of the Theoretical High Energy Physics Group at Institut Teknologi Bandung for the discussions, insights, and hospitality.
Appendix A The expression of $e^{2\nu}$ for the isotropic configuration with the limit $\mathcal{B}\rightarrow 0$
For the isotropic non-magnetized configuration,
$\alpha=K,\>\>\>\>\beta=1,\>\>\>\>\gamma=0.$ (82)
In this case, Eq. (LABEL:numagnetized) gives
$\nu$ $=$ $\mathscr{C}-(2-1)\bigg[\ln\{K\rho^{2}+\rho\}$ (83)
$-\ln\bigg\{\frac{2K\rho+1-1}{2K\rho+1+1}\bigg\}\bigg].$
Here, $\mathscr{C}$ is a constant. Further calculation gives
$\nu=$ $\mathscr{C}-\ln[\rho(K\rho+1)]+\ln\bigg[\frac{K\rho}{K\rho+1}\bigg].$
$\nu=$ $\mathscr{C}+\ln K-\ln(K\rho+1)-\ln(K\rho+1).$ (84)
We can define $\nu_{c}\equiv\mathscr{C}+\ln K$, so that
$\nu=\nu_{c}-2\ln(K\rho+1).$ (85)
The metric function $e^{2\nu(r)}$ now reads
$e^{2\nu}=\frac{e^{2\nu_{c}}}{\bigg[1+K\rho_{c}\left\{1-\frac{(r-r_{0})^{2}}{(R_{s}-r_{0})^{2}}\right\}\bigg]^{4}}.$ (86)
Eq. (86) is nothing but Eq. (24) in Ref. Nojiri2024 .
Now we glance at the toy model started from Eq. (79). For the non-magnetized configuration, the mathematical expression will be
$\nu$ $=$ $\mathscr{C}-\ln\!\big[\rho(K\rho+1)\big]$ (87)
$+\frac{1}{\sqrt{1}}\ln\!\left(\frac{2K\rho+1-1}{2K\rho+1+1}\right).$
Note that Eq. (87) is mathematically identical to Eq. (83), which implies that the metric function of the toy model shown by Eq. (80) can also be reduced to Eq. (86).
Appendix B The expression of $(e^{2\nu})^{\prime}$ for the isotropic configuration with the limit $\mathcal{B}\rightarrow 0$
Note that the condition shown by Eq. (82) is also satisfied in this case. Eq. (LABEL:derivativenu) is now reduced and becomes
$\left(e^{2\nu}\right)^{\prime}$ $=$ $e^{2\nu}\,4(2-1)\,\frac{\rho_{c}(r-r_{0})}{(R_{s}-r_{0})^{2}}$ (88)
$\times\left(\frac{2K\rho+1}{K\rho^{2}+\rho}-\frac{1}{K\rho^{2}+\rho}\right)$
$=$ $\frac{8K\rho_{c}(r-r_{0})(R_{s}-r_{0})^{-2}}{K\rho+1}$
$\times\frac{e^{2\nu_{c}}}{(1+K\rho)^{4}}.$
Finally, Eq. (88) can be written as
$(e^{2\nu})^{\prime}=\frac{8K\rho_{c}e^{2\nu_{c}}(r-r_{0})(R_{s}-r_{0})^{-2}}{\bigg[1+K\rho_{c}\left\{1-\frac{(r-r_{0})^{2}}{(R_{s}-r_{0})^{2}}\right\}\bigg]^{5}},$ (89)
which is actually Eq. (24) in Ref. Nojiri2024 .